OpenStudy (anonymous):

A wheel has a diameter of 16 cm. Approximately how far does it travel in 5 revolutions? A. 502.4 cm B. 251.2 cm C. 100.48 cm D. 80 cm

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@IMStuck

OpenStudy (dejmccray):

ok do C=d*3.14 so 16*3.14=50.24 then multiply that by 5 it = 251.20 which would be B

OpenStudy (imstuck):

Ok, here's how you do this. If the circle has a diameter, what is its circumference? \[C = 2\pi r\]or \[C=\pi d\]Your diameter is 16. So your formula is \[C=16 \pi \]So in 5 full revolutions you go around 5 times 16 pi, or 80 pi. 80 times pi = 251.3 or B